the perimiter of a rectangle garden is 136feet the length of the garden is 4 feet less than 3 times the width write a system of equations that can be used to find the measure of the length and width of the rectangle let y be the measure of length and x be the measure of width

Answer by jorel1380(3719)

If X is the measure of width and Y the length, then our equations are:
Width=X
Length=Y=3X-4
Since our perimeter is 136 ft.:
2(X)+2(3X-4)=136
2X+6X-8=136
8X=144
X=18
Y=3X-4=50